OpenStudy (anonymous):

Find the next three terms of the sequence 4, 5, 9, 14, 23, 37

OpenStudy (zzr0ck3r):

60,97,157

OpenStudy (anonymous):

Thank You how are you getting that?

OpenStudy (zzr0ck3r):

5+4 = 9 14+9 = 23 23+14=37 .......

OpenStudy (anonymous):

okay thank you

OpenStudy (zzr0ck3r):

np